# Spectral Convolutional Conditional Neural Processes

### v5

@@ 3 Spectral Convolutional Conditional Neural Processes
The decoder $\varphi_{d}$ in ConvCNPs is typically parameterized using standard CNNs such as U-Net (ronneberger2015u) or ResNet (he2016deep). These architectures employ discrete convolutional kernels—finite sets of learnable parameters that define localized filters operating over neighboring grid points. The kernel size, fixed a priori, determines the receptive field of each convolution (ding2022scaling) and is generally much smaller than the overall extent of the input signals in the physical domain (romero2021ckconv; knigge2023modelling). This locality constraint fundamentally limits a model’s ability to capture long-range dependencies and to integrate information from observations distributed across large spatial or temporal domains (Peng_2017_CVPR; wang2018non; ramachandran2019stand; wang2020axial). The issue becomes particularly pronounced when handling sparse or irregularly sampled data, where effective global reasoning cannot emerge solely from local convolution operations. Although enlarging the convolutional kernel increases the receptive field, it results in a rapid escalation of both parameter count and computational cost. Transformer-based architectures mitigate this issue by enabling explicit global interactions (vaswani2017attention); however, they incur quadratic, rather than linear, complexity in the number of inputs, rendering them impractical for large context sets unless one resorts to approximation schemes (nguyen2022transformer; feng2022latent; ashman2024translation; ashman2025gridded).
@@
The computational cost of SConvCNPs comprises three parts: (i) $\mathcal{O}(|\mathcal{D}_{c}||\mathcal{G}|)$ to compute the discretized functional embedding on the grid $\mathcal{G}$ (equation 3); (ii) $\mathcal{O}(|\mathcal{G}|\log|\mathcal{G}|)$ for the FFT-based spectral convolutions (equation 1); and (iii) $\mathcal{O}(|\mathcal{D}_{q}||\mathcal{G}|)$ to interpolate grid embeddings at the query locations (equation 4). Overall, the total complexity is $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+\log|\mathcal{G}|+|\mathcal{D}_{q}|)\bigr)$, matching the $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+1+|\mathcal{D}_{q}|)\bigr)$ complexity of ConvCNPs up to logarithmic factors. Both architectures therefore scale linearly with task size. A key limitation, however, is that $|\mathcal{G}|$ grows exponentially with the input dimension $d_{x}$, restricting these methods to low-dimensional domains. In contrast, transformer-based NPs (TNPs; kim2019attentive; nguyen2022transformer; feng2022latent; ashman2024translation) avoid gridding entirely and thus scale more gracefully with input dimensionality. Instead, their computational cost scales quadratically with the task size, typically as $\mathcal{O}(|\mathcal{D}_{c}|^{2}+|\mathcal{D}_{q}|^{2})$, with minor variations depending on the specific implementation (see Section B.1.1).
@@
The convolution operator preserves translation equivariance under the Fourier transform (see equation 1). However, practical FNOs implementations often include explicit positional information to improve predictive accuracy (li2020fourier; tran2021factorized; gupta2021multiwavelet; rahman2022uno; helwig2023group; tripura2023wavelet; liu2023domain; li2024multi). Accordingly, we augment the functional embedding $\varphi_{e}[\mathcal{D}_{c}](x)$ with positional information:
@@
While this augmentation explicitly breaks translation equivariance, we empirically observe performance improvements (see Section B.5). An interesting direction for future work is to investigate relative positional encodings (shaw2018self; su2024roformer), which can provide spatial context while preserving translation equivariance.
@@
Unlike the continuous Fourier transform, the DFT—and therefore the FFT—depends inherently on the grid $\mathcal{G}$ on which $\varphi_{e}[\mathcal{D}_{c}]$ (or $\widetilde{\varphi}_{e}[\mathcal{D}_{c}]$) is represented. This dependence arises from both the grid resolution and its physical extent: changing either alters the resulting Fourier coefficients and can lead to inconsistent behavior in the outputs (see Section A). Sensitivity to resolution is not unique to DFT; CNNs exhibit analogous issues (raonic2023convolutional; bartolucci2023representation). For example, ConvCNPs mitigate this effect by fixing the grid resolution. While spatial CNNs become stable once the resolution is fixed, DFT-based methods remain sensitive unless both the resolution and the physical range are controlled. Accordingly, we fix both, choosing a domain sufficiently large to cover all context and query inputs across tasks. When this is impractical, the domain can instead be divided into (possibly overlapping) fixed-size patches, with spectral convolutions applied independently to each patch and the outputs aggregated—mirroring the mechanism of standard convolution layers. Related ideas in efficient transformer architectures suggest that this is a promising direction for future work (beltagy2020longformer; zaheer2020big; liu2021swin; ding2023longnet).
@@ 4 Experiments
We evaluate our framework on four regression benchmarks and compare its performance against several members of CNPs family. Specifically, we include the original CNP (garnelo2018conditional), the Attentive CNP (AttCNP; kim2019attentive), the Convolutional CNP (ConvCNP; gordon2019convolutional), the diagonal variant of the Transformer Neural Process (TNP; nguyen2022transformer), and the Translation-Equivariant Transformer Neural Process (TETNP; ashman2024translation). For each experimental setting, all models are trained using four random seeds. Performance is evaluated using the final aggregated log-likelihood and root-mean-squared error (RMSE), and we report the mean $\pm$ standard deviation across runs. Additional details regarding datasets, architectures, and training procedures are provided in Appendix B. Our implementation and experimental code are publicly available at https://github.com/peiman-m/SConvCNP.
@@ 4.1 Synthetic 1-D Regression
We begin by evaluating models on four synthetic benchmarks generated from distinct stochastic processes: a GP with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. For each benchmark, the parameters of the generative process—kernel hyperparameters for the GPs, frequency and direction for the sawtooth generator, and frequency and duty cycle for the square-wave generator—are sampled randomly. Complete experimental details are provided in Appendix B.1. Table 1 reports the average evaluation metrics over 1,000 test batches, each containing 16 tasks (each task defined in equation 2). For every task—independent of batch—both $x_{c}$ and $x_{q}$ are sampled independently from $\mathcal{U}[-3,3)$. For each test batch, the number of context points, shared by all tasks within that batch, is drawn independently according to $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$. The number of query points, however, is fixed at $|\mathcal{I}_{q}|=256$ for all the test tasks. As shown, SConvCNP consistently outperforms or closely matches strong baselines, including TETNP and ConvCNP, which represent the current state of the art. Figure 1 provides qualitative comparisons of predictive maps produced by SConvCNP, ConvCNP, and TETNP.
For the sawtooth-wave benchmark, we were unable to successfully train either TNP or TETNP. Across the configurations we attempted—including increased model capacity and multiple randomized initializations—their predictions consistently collapsed to zero. We hypothesize that this failure mode stems from spectral bias, the tendency of neural networks to favor low-frequency structure over high-frequency components (rahaman2019spectral; ronen2019convergence; basri2020frequency; tancik2020fourier; fridovich2022spectral). Recent findings by vasudeva2025transformers further suggest that this bias is exacerbated in transformer architectures compared with convolutional networks. This interpretation aligns with observations by nguyen2022transformer, who originally introduced TNPs. They report degraded performance on GP samples drawn from a periodic kernel—reflected in poor log-likelihood scores—despite strong results on Matérn-kernel tasks. Interestingly, our experiments on periodic-kernel GP tasks (Table 1) do not replicate this limitation. A key distinction is that we employ the Efficient Query TNP (EQTNP) architecture of feng2022latent, rather than the original TNP design (nguyen2022transformer) (see Appendix B.1.1 for architectural details). Finally, although sawtooth and square-wave functions share similar discontinuity and high-frequency characteristics, Table 1 shows that TNP and TETNP do not collapse when trained on square-wave signals. Investigating the source of this discrepancy falls beyond the scope of this work but represents an intriguing direction for future study.
@@ 4.2 Predator–Prey Dynamics
We next assess performance on simulated trajectories from a stochastic variant of the Lotka–Volterra predator–prey system (Lotka1910contribution; Volterra1926variazioni), following the formulation in bruinsma2023autoregressive. Let $U_{t}$ and $V_{t}$ denote the prey and predator populations at time $t$, respectively. Their dynamics evolve according to the stochastic Lotka–Volterra system
@@
Table 2 reports average evaluation metrics over 1,000 test batches, each containing 16 tasks. For each batch, the number of context points—shared across all tasks—is sampled as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$, and the number of query points is fixed at $|\mathcal{I}_{q}|=256$. SConvCNP achieves log-likelihood performance comparable to ConvCNP, and both are competitive with TETNP. In terms of RMSE, SConvCNP matches TETNP, indicating that its predictive uncertainty intervals are slightly wider. Qualitative comparisons of predictive maps for the top three models appear in Figure 2.
@@ 4.3 Traffic Flow
For our third experiment, we evaluate on the California traffic-flow dataset from LargeST (liu2023largest). This dataset comprises five years (2017–2021) of traffic measurements recorded every 5 minutes by approximately 8,600 loop-detector sensors deployed across California’s highway network. We focus on the year 2020, where traffic patterns are expected to exhibit heightened variability due to the abrupt onset of the COVID-19 pandemic. For each sensor, we segment its year-long time series into non-overlapping 14-day windows and downsample each window by a factor of 6 (from 5-minute to 30-minute resolution). Each window is treated as a dense trajectory from which individual tasks are constructed. Additional details of the experimental setup are provided in Appendix B.3.
Table 3 reports evaluation metrics averaged over 66,586 test tasks, constructed from 2,561 held-out test sensors and partitioned into batches of size 32. For each batch, the number of context points—shared across all tasks—is drawn as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$, and the number of query points is fixed at $|\mathcal{I}_{q}|=50$. Across all metrics, SConvCNP achieves the best performance, attaining the highest log-likelihoods and lowest RMSE among the baselines. Figure 3 provides qualitative comparisons of the predictive maps produced by the three models.
@@ 4.4 Image Completion
For our final experiment, we evaluate model performance on an image-completion task formulated as a spatial regression problem, where the model maps 2D pixel coordinates to their corresponding intensity values. We use images from the Describable Textures Dataset (DTD; cimpoi2014describing), and construct each task from a processed $64\times 64$ subsampled crop of an original image. Because each task contains a large number of context and query pixels, we were unable to fit TETNP within our computational budget—even after reducing its size—so we exclude it from this experiment. Additional experimental details are provided in Appendix B.4. Table 4 reports results averaged over 1,880 test tasks, evaluated in batches of 16. For each batch, the number of context points—shared across all tasks—is sampled as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,1024)$; all remaining pixels serve as query points, so that $|\mathcal{I}_{q}|=64\times 64-|\mathcal{I}_{c}|$. Among the baselines, SConvCNP achieves the highest log-likelihoods and the second-lowest RMSE. Figure 4 presents qualitative comparisons of the predictive means produced by the top three models.
@@ 5 Related Works
The high computational cost of traditional numerical PDE solvers has motivated the development of more efficient alternatives based on deep learning gupta2022towards. Among these, neural operators (li2020neural; kovachki2023neural), and in particular Fourier neural operators (FNOs, li2020fourier), have become popular. Numerous extensions have since been proposed: helwig2023group introduced rotation- and reflection-equivariant FNOs; gupta2021multiwavelet developed multiwavelet neural operators by projecting kernels onto predefined polynomial bases; and tran2021factorized reduced model complexity through separable Fourier representations. xiao2024amortized proposed the Amortized FNO (AM-FNO), which substantially cuts parameter count by amortizing the Fourier kernel parameterization, while qin2024toward analyzed FNOs through the lens of spectral bias. koshizuka2024understanding provided a mean-field theoretical perspective, and zheng2024alias introduced the Mamba Neural Operator, which couples global context via a state-space model to achieve linear complexity and representation equivalence. bartolucci2023representation examined continuous–discrete consistency, showing that it holds for convolutional neural operators (raonic2023convolutional). Additional works include orthogonal-attention eigenfunction methods for operator learning (xiao2023improved) and hierarchical transformers with frequency-aware priors for resolution-invariant super-resolution (luo2024hierarchical).
@@
In non-parametric Bayesian modeling, GPs and deep GPs (damianou2013deep) provide flexible function-space priors with well-calibrated uncertainty estimates, but their computational cost becomes prohibitive on large datasets. This challenge has motivated the development of Bayesian neural networks (BNNs; mackay1992practical; hinton1993keeping; neal2012bayesian), which combine neural network expressiveness with Bayesian uncertainty quantification. However, specifying meaningful priors over network weights remains notoriously difficult. Recent work therefore reframes Bayesian inference in neural networks as learning a posterior over the functions induced by stochastic weights (wolpert1993bayesian; qiu2023should). Variational implicit processes (VIPs; ma2019variational; santana2021function; ortega2022deep) generalize GPs by defining implicit stochastic processes through latent variables, while functional variational BNNs (fBNNs; sun2019functional) enforce alignment between BNN-induced priors and target priors by minimizing a functional KL divergence—though this objective can be difficult to compute and, in some cases, ill-posed (burt2020understanding). Subsequent work has aimed to address these limitations (ma2021functional; rudner2022tractable; wild2022generalized; rudner2023function; wu2024functional). Orthogonally, energy-based models have been explored for representing stochastic processes in function space (yang2020energy; lim2022energy). More recently, several approaches have extended diffusion and flow models to function spaces (phillips2022spectral; dutordoir2023neural; mathieu2023geometric; franzese2023continuous; lim2023score; pmlr-v238-kerrigan24a; zhang2025flow).

### v6

@@ 3 Spectral Convolutional Conditional Neural Processes
The decoder $\varphi_{d}$ in ConvCNPs is typically parameterized using standard CNNs such as U-Net (Ronneberger et al., 2015) or ResNet (He et al., 2016). These architectures employ discrete convolutional kernels—finite sets of learnable parameters that define localized filters operating over neighboring grid points. The kernel size, fixed a priori, determines the receptive field of each convolution (Ding et al., 2022) and is generally much smaller than the overall extent of the input signals in the physical domain (Romero et al., 2021; Knigge et al., 2023). This locality constraint fundamentally limits a model’s ability to capture long-range dependencies and to integrate information from observations distributed across large spatial or temporal domains (Peng et al., 2017; Wang et al., 2018; Ramachandran et al., 2019; Wang et al., 2020). The issue becomes particularly pronounced when handling sparse or irregularly sampled data, where effective global reasoning cannot emerge solely from local convolution operations. Although enlarging the convolutional kernel increases the receptive field, it results in a rapid escalation of both parameter count and computational cost. Transformer-based architectures mitigate this issue by enabling explicit global interactions (Vaswani et al., 2017); however, they incur quadratic, rather than linear, complexity in the number of inputs, rendering them impractical for large context sets unless one resorts to approximation schemes (Nguyen and Grover, 2022; Feng et al., 2022; Ashman et al., 2024a, 2025).
@@
The computational cost of SConvCNPs comprises three parts: (i) $\mathcal{O}(|\mathcal{D}_{c}||\mathcal{G}|)$ to compute the discretized functional embedding on the grid $\mathcal{G}$ (equation 3); (ii) $\mathcal{O}(|\mathcal{G}|\log|\mathcal{G}|)$ for the FFT-based spectral convolutions (equation 1); and (iii) $\mathcal{O}(|\mathcal{D}_{q}||\mathcal{G}|)$ to interpolate grid embeddings at the query locations (equation 4). Overall, the total complexity is $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+\log|\mathcal{G}|+|\mathcal{D}_{q}|)\bigr)$, matching the $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+1+|\mathcal{D}_{q}|)\bigr)$ complexity of ConvCNPs up to logarithmic factors. Both architectures therefore scale linearly with task size. A key limitation, however, is that $|\mathcal{G}|$ grows exponentially with the input dimension $d_{x}$, restricting these methods to low-dimensional domains. In contrast, transformer-based NPs (TNPs; Kim et al. (2019); Nguyen and Grover (2022); Feng et al. (2022); Ashman et al. (2024a)) avoid gridding entirely and thus scale more gracefully with input dimensionality. Instead, their computational cost scales quadratically with the task size, typically as $\mathcal{O}(|\mathcal{D}_{c}|^{2}+|\mathcal{D}_{q}|^{2})$, with minor variations depending on the specific implementation (see Section B.1.1).
@@
The convolution operator preserves translation equivariance under the Fourier transform (see equation 1). However, practical FNOs implementations often include explicit positional information to improve predictive accuracy (Li et al., 2020a; Tran et al., 2021; Gupta et al., 2021; Rahman et al., 2022b; Helwig et al., 2023; Tripura and Chakraborty, 2023; Liu et al., 2023a; Li et al., 2024). Accordingly, we augment the functional embedding $\varphi_{e}[\mathcal{D}_{c}](x)$ with positional information:
@@
While this augmentation explicitly breaks translation equivariance, we empirically observe performance improvements (see Section B.5). An interesting direction for future work is to investigate relative positional encodings (Shaw et al., 2018; Su et al., 2024), which can provide spatial context while preserving translation equivariance.
@@
Unlike the continuous Fourier transform, the DFT—and therefore the FFT—depends inherently on the grid $\mathcal{G}$ on which $\varphi_{e}[\mathcal{D}_{c}]$ (or $\widetilde{\varphi}_{e}[\mathcal{D}_{c}]$) is represented. This dependence arises from both the grid resolution and its physical extent: changing either alters the resulting Fourier coefficients and can lead to inconsistent behavior in the outputs (see Section A). Sensitivity to resolution is not unique to DFT; CNNs exhibit analogous issues (Raonic et al., 2023; Bartolucci et al., 2023). For example, ConvCNPs mitigate this effect by fixing the grid resolution. While spatial CNNs become stable once the resolution is fixed, DFT-based methods remain sensitive unless both the resolution and the physical range are controlled. Accordingly, we fix both, choosing a domain sufficiently large to cover all context and query inputs across tasks. When this is impractical, the domain can instead be divided into (possibly overlapping) fixed-size patches, with spectral convolutions applied independently to each patch and the outputs aggregated—mirroring the mechanism of standard convolution layers. Related ideas in efficient transformer architectures suggest that this is a promising direction for future work (Beltagy et al., 2020; Zaheer et al., 2020; Liu et al., 2021; Ding et al., 2023).
@@ 4 Experiments
We evaluate our framework on four regression benchmarks and compare its performance against several members of CNPs family. Specifically, we include the original CNP (Garnelo et al., 2018a), the Attentive CNP (AttCNP; Kim et al. (2019)), the Convolutional CNP (ConvCNP; Gordon et al. (2019)), the diagonal variant of the Transformer Neural Process (TNP; Nguyen and Grover (2022)), and the Translation-Equivariant Transformer Neural Process (TETNP; Ashman et al. (2024a)). For each experimental setting, all models are trained using four random seeds. Performance is evaluated using the final aggregated log-likelihood and root-mean-squared error (RMSE), and we report the mean $\pm$ standard deviation across runs. Additional details regarding datasets, architectures, and training procedures are provided in Appendix B. Our implementation and experimental code are publicly available at https://github.com/peiman-m/SConvCNP.
@@ 4.1 Synthetic 1-D Regression
We begin by evaluating models on four synthetic benchmarks generated from distinct stochastic processes: a GP with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. For each benchmark, the parameters of the generative process—kernel hyperparameters for the GPs, frequency and direction for the sawtooth generator, and frequency and duty cycle for the square-wave generator—are sampled randomly. Complete experimental details are provided in Appendix B.1. Table 1 reports the average evaluation metrics over 1,000 test batches, each containing 16 tasks (each task defined in equation 2). For every task—independent of batch—both $x_{c}$ and $x_{q}$ are sampled independently from $\mathcal{U}[-3,3)$. For each test batch, the number of context points, shared by all tasks within that batch, is drawn independently according to $n_{c}\sim\mathcal{U}[5,25)$. The number of query points, however, is fixed at $n_{q}=256$ for all the test tasks. As shown, SConvCNP consistently outperforms or closely matches strong baselines, including TETNP and ConvCNP, which represent the current state of the art. Figure 1 provides qualitative comparisons of predictive maps produced by SConvCNP, ConvCNP, and TETNP.
For the sawtooth-wave benchmark, we were unable to successfully train either TNP or TETNP. Across the configurations we attempted—including increased model capacity and multiple randomized initializations—their predictions consistently collapsed to zero. We hypothesize that this failure mode stems from spectral bias, the tendency of neural networks to favor low-frequency structure over high-frequency components (Rahaman et al., 2019; Ronen et al., 2019; Basri et al., 2020; Tancik et al., 2020; Fridovich-Keil et al., 2022). Recent findings by Vasudeva et al. (2025) further suggest that this bias is exacerbated in transformer architectures compared with convolutional networks. This interpretation aligns with observations by Nguyen and Grover (2022), who originally introduced TNPs. They report degraded performance on GP samples drawn from a periodic kernel—reflected in poor log-likelihood scores—despite strong results on Matérn-kernel tasks. Interestingly, our experiments on periodic-kernel GP tasks (Table 1) do not replicate this limitation. A key distinction is that we employ the Efficient Query TNP (EQTNP) architecture of Feng et al. (2022), rather than the original TNP design (Nguyen and Grover, 2022) (see Appendix B.1.1 for architectural details). Finally, although sawtooth and square-wave functions share similar discontinuity and high-frequency characteristics, Table 1 shows that TNP and TETNP do not collapse when trained on square-wave signals. Investigating the source of this discrepancy falls beyond the scope of this work but represents an intriguing direction for future study.
@@ 4.2 Predator–Prey Dynamics
We next assess performance on simulated trajectories from a stochastic variant of the Lotka–Volterra predator–prey system (Lotka, 1910; Volterra, 1926), following the formulation in Bruinsma et al. (2023). Let $U_{t}$ and $V_{t}$ denote the prey and predator populations at time $t$, respectively. Their dynamics evolve according to the stochastic Lotka–Volterra system
@@
Table 2 reports average evaluation metrics over 1,000 test batches, each containing 16 tasks. For each batch, the number of context points—shared across all tasks—is sampled as $n_{c}\sim\mathcal{U}[5,25)$, and the number of query points is fixed at $n_{q}=256$. SConvCNP achieves log-likelihood performance comparable to ConvCNP, and both are competitive with TETNP. In terms of RMSE, SConvCNP matches TETNP, indicating that its predictive uncertainty intervals are slightly wider. Qualitative comparisons of predictive maps for the top three models appear in Figure 2.
@@ 4.3 Traffic Flow
For our third experiment, we evaluate on the California traffic-flow dataset from LargeST (Liu et al., 2023b). This dataset comprises five years (2017–2021) of traffic measurements recorded every 5 minutes by approximately 8,600 loop-detector sensors deployed across California’s highway network. We focus on the year 2020, where traffic patterns are expected to exhibit heightened variability due to the abrupt onset of the COVID-19 pandemic. For each sensor, we segment its year-long time series into non-overlapping 14-day windows and downsample each window by a factor of 6 (from 5-minute to 30-minute resolution). Each window is treated as a dense trajectory from which individual tasks are constructed. Additional details of the experimental setup are provided in Appendix B.3.
Table 3 reports evaluation metrics averaged over 66,586 test tasks, constructed from 2,561 held-out test sensors and partitioned into batches of size 32. For each batch, the number of context points—shared across all tasks—is drawn as $n_{c}\sim\mathcal{U}[5,25)$, and the number of query points is fixed at $n_{q}=50$. Across all metrics, SConvCNP achieves the best performance, attaining the highest log-likelihood and lowest RMSE among the baselines. Figure 3 provides qualitative comparisons of the predictive maps produced by the three models.
@@ 4.4 Image Completion
For our final experiment, we evaluate model performance on an image-completion task formulated as a spatial regression problem, where the model maps 2D pixel coordinates to their corresponding intensity values. We use images from the Describable Textures Dataset (DTD; Cimpoi et al. (2014)), and construct each task from a processed $64\times 64$ subsampled crop of an original image. Because each task contains a large number of context and query pixels, we were unable to fit TETNP within our computational budget—even after reducing its size—so we exclude it from this experiment. Additional experimental details are provided in Appendix B.4. Table 4 reports results averaged over 1,880 test tasks, evaluated in batches of 16. For each batch, the number of context points—shared across all tasks—is sampled as $n_{c}\sim\mathcal{U}[5,1024)$; all remaining pixels serve as query points, so that $n_{q}=4096-n_{c}$. Among the baselines, SConvCNP achieves the highest log-likelihood and the second-lowest RMSE. Figure 4 presents qualitative comparisons of the predictive means produced by the top three models.
@@ 5 Related Works
The high computational cost of traditional numerical PDE solvers has motivated the development of more efficient alternatives based on deep learning Gupta and Brandstetter (2022). Among these, neural operators (Li et al., 2020b; Kovachki et al., 2023), and in particular Fourier neural operators (FNOs, Li et al., 2020a), have become popular. Numerous extensions have since been proposed: Helwig et al. (2023) introduced rotation- and reflection-equivariant FNOs; Gupta et al. (2021) developed multiwavelet neural operators by projecting kernels onto predefined polynomial bases; and Tran et al. (2021) reduced model complexity through separable Fourier representations. Xiao et al. (2024) proposed the Amortized FNO (AM-FNO), which substantially cuts parameter count by amortizing the Fourier kernel parameterization, while Qin et al. (2024) analyzed FNOs through the lens of spectral bias. Koshizuka et al. (2024) provided a mean-field theoretical perspective, and Zheng et al. (2024) introduced the Mamba Neural Operator, which couples global context via a state-space model to achieve linear complexity and representation equivalence. Bartolucci et al. (2023) examined continuous–discrete consistency, showing that it holds for convolutional neural operators (Raonic et al., 2023). Additional works include orthogonal-attention eigenfunction methods for operator learning (Xiao et al., 2023) and hierarchical transformers with frequency-aware priors for resolution-invariant super-resolution (Luo et al., 2024).
@@
In non-parametric Bayesian modeling, GPs and deep GPs (Damianou and Lawrence, 2013) provide flexible function-space priors with well-calibrated uncertainty estimates, but their computational cost becomes prohibitive on large datasets. This challenge has motivated the development of Bayesian neural networks (BNNs; MacKay (1992); Hinton and Van Camp (1993); Neal (2012)), which combine neural network expressiveness with Bayesian uncertainty quantification. However, specifying meaningful priors over network weights remains notoriously difficult. Recent work therefore reframes Bayesian inference in neural networks as learning a posterior over the functions induced by stochastic weights (Wolpert, 1993; Qiu et al., 2023). Variational implicit processes (VIPs; Ma et al. (2019); Santana et al. (2021); Ortega et al. (2022)) generalize GPs by defining implicit stochastic processes through latent variables, while functional variational BNNs (fBNNs; Sun et al. (2019)) enforce alignment between BNN-induced priors and target priors by minimizing a functional KL divergence—though this objective can be difficult to compute and, in some cases, ill-posed (Burt et al., 2020). Subsequent work has aimed to address these limitations (Ma and Hernández-Lobato, 2021; Rudner et al., 2022; Wild et al., 2022; Rudner et al., 2023; Wu et al., 2024). Orthogonally, energy-based models have been explored for representing stochastic processes in function space (Yang et al., 2020; Lim et al., 2022). More recently, several approaches have extended diffusion and flow models to function spaces (Phillips et al., 2022; Dutordoir et al., 2023; Mathieu et al., 2023; Franzese et al., 2023; Lim et al., 2023; Kerrigan et al., 2024; Zhang and Scott, 2025).